Two angles form a linear pair. The measure of one angle is 1/3 the measure of the other angle.
Answer by stanbon(75887)

Two angles form a linear pair. The measure of one angle is 1/3 the measure of the other angle.
------
Equations:
x + y = 180 degrees
x = (1/3)y
-----
Substitute for "x" and solve for "y"::
(1/3)y + y = 180
(4/3)y = 180
y = (3/4)180
y = 135 degrees
----
Solve for "x":
x + y = 180
x + 135 = 180
x = 45 degrees
---------
Cheers,
Stan H.
-----------------